# Split and Rephrase with Large Language Models

### v2

@@ Appendix B Split and Rephrase Examples
In Tables 6, 7 and 8, we provide illustrative examples of split and rephrased output obtained with the fine-funed Pythia-6.9B model on Desse, BiSECT and Wiki-BM source sentences, respectively. Sequences from the source text that have been correctly deleted or modified are shown in purple. Sequences that have been correctly added to the output are shown in green. Sequences that have been incorrectly altered in the original sentence are shown in red.
@@
Most people travel to Europe or the Islands and get a taste of the culture in those places.
Most people travel to Europe or the Islands. They get a taste of the culture in those places.
People would’ve been killed, places would have been bombed and there would have been so much going on all at once.
People would’ve been killed. Places would have been bombed. There would have been so much going on all at once.
War is injuring innocent people which is not necessary.
War is injuring innocent people. It is not necessary.
As education continues to reach the poorer areas of our world i truly believe that one day we will all live in TRUE harmony.
Education continues to reach the poorer areas of our world. I truly believe that one day we will all live in TRUE harmony.
This land is still the home to many civilians who rely on the resources in their area to live.
This land is still the home to many civilians. They rely on the resources in their area to live.
@@
Dialysis does not alter serum doxycycline half-life and thus would not be of benefit in treating cases of overdose.
Dialysis does not change the serum half-life of doxycycline. Thus, it would not be of benefit in the treatment of cases of overdose.
In the first study, type A resistant patients (A-non responders) were randomised to receive placebo or 10, 000 U of NeuroBloc and in the second, type A toxin responsive patients (A-responders) were randomised to receive placebo, 5000 U or 10, 000 U of toxin.
In the first study, type A non-responders were randomized to receive placebo or 10, 000 U of NeuroBloc. In the second study, type A responders were randomized to receive placebo, 5000 U or 10, 000 U of toxin.
Most patients treated with Zevalin experience a severely reduced number of blood cells, which needs some time to normalise again after treatment.
Most patients treated with Zevalin experience a severe reduction in the number of blood cells. It takes some time to return to normal after treatment.
Lisonorm tablet contains 5mg amlodipine/ 10mg lisinopril and is an anti-hypertensive indicated in patients with blood pressure adequately controlled with lisinopril and amlodipine given concurrently at the same dose level.
Lisonorm tablet contains 5 mg of amlodipine and 10 mg of lisinopril. It is indicated for the treatment of hypertension in patients whose blood pressure is adequately controlled with lisinopril and amlodipine.
@@
She walked out on their wedding , and she currently is single and holds intense anger towards Richard for ruining their wedding, and also became furious to learn she is pregnant, even though Richard told her he got a vasectomy.
She walked out on their wedding. She currently is single. She holds intense anger towards Richard for ruining their wedding. She became furious to learn she is pregnant. Richard told her he got a vasectomy.
Southfort is an unincorporated community in DeSoto County, Florida, United States, located approximately south of Fort Ogden on U.S. Route 17.
Southfort is an unincorporated community in DeSoto County, Florida. It is located approximately south of Fort Ogden on U.S. Route 17.
Parke played for powerhouse youth club FC Delco, and played four years of college soccer at Drexel University, before being drafted by the MetroStars with 60th and last overall pick of the 2004 MLS SuperDraft.
Parke played for powerhouse youth club FC Delco. He played four years of college soccer at Drexel University. He was drafted by the MetroStars with 60th and last overall pick of the 2004 MLS SuperDraft.
Auguste Le Duc started on a bridge between the two islands, which was quickly swept away by the forces of nature.
Auguste Le Duc started on a bridge between the two islands. The bridge was quickly swept away by the forces of nature.
@@
Appendix C Evaluation Metrics on Wiki-BM
@@
In section 7, we used portions of the Wiki-BM corpus (Zhang et al., 2020a) for our qualitative evaluation of LLM-based SPRP output. In Table 9, we provide the results in terms of metrics for the complete Wiki-BM test set, along with results from AlAjlouni and Li (2023), as representative of an alternative method based on fine-tuning an encoder-decoder pretrained T5 model.
@@
The results for our LLM-based approach are in line with those obtained on the DeSSE test set across metrics: slightly lower on BESC, BLEU, and SARI, but slightly higher in terms of compliance (CPL). Our model outperformed the reported scores in AlAjlouni and Li (2023) by 13.4 and 5.1 points on the BLEU and SARI metrics, respectively.
@@
pythia-6.9B-DE 72.9 65.1 36.14 91.15 68.10 85.50 98.19
[…]
Table 9: Metrics results on the Wiki-BM benchmark using the Pythia-6.9B model fine-tuned on DeSSE data. Available results from alternative approaches in terms of BLEU and SARI are reported as per the cited work.
Appendix D 3-Way Ranking Evaluation Protocol
@@
The 3-way ranking manual evaluation involved 11 human evaluators, all researchers with at least upper-intermediate knowledge of English, who were not paid for the task. The evaluation was split into two tasks: for one task, the evaluators were asked to rank the output of the fine-tuned Pythia-6.9B model and the output of the ABCD model; for the other task, they had to rank the output of the fine-tuned Pythia-6.9B model and the output of the BiSECT model. Each of the two tasks involved 100 samples.
@@
Appendix E Qualitative Evaluation Protocol
@@
Appendix F Initial Multilingual Results

### v4

@@ Appendix B Split and Rephrase Examples
In Tables 7, 8 and 9, we provide illustrative examples of split and rephrased output obtained with the fine-funed Pythia-6.9B model on Desse, BiSECT and Wiki-BM source sentences, respectively. We selected examples that are representative of the type of alterations to the source performed by the selected model, which was used for the manual evaluations, including errors in splitting and/or rephrasing.
Sequences from the source text that have been correctly deleted or modified are shown in purple. Sequences that have been correctly added to the output are shown in green. Sequences that have been incorrectly altered in the original sentence are shown in red.
@@
Most people travel to Europe or the Islands and get a taste of the culture in those places. Most people travel to Europe or the Islands. They get a taste of the culture in those places.
People would’ve been killed, places would have been bombed and there would have been so much going on all at once. People would’ve been killed. Places would have been bombed. There would have been so much going on all at once.
War is injuring innocent people which is not necessary. War is injuring innocent people. It is not necessary.
As education continues to reach the poorer areas of our world i truly believe that one day we will all live in TRUE harmony. Education continues to reach the poorer areas of our world. I truly believe that one day we will all live in TRUE harmony.
This land is still the home to many civilians who rely on the resources in their area to live. This land is still the home to many civilians. They rely on the resources in their area to live.
@@
Dialysis does not alter serum doxycycline half-life and thus would not be of benefit in treating cases of overdose. Dialysis does not change the serum half-life of doxycycline. Thus, it would not be of benefit in the treatment of cases of overdose.
In the first study, type A resistant patients (A-non responders) were randomised to receive placebo or 10, 000 U of NeuroBloc and in the second, type A toxin responsive patients (A-responders) were randomised to receive placebo, 5000 U or 10, 000 U of toxin. In the first study, type A non-responders were randomized to receive placebo or 10, 000 U of NeuroBloc. In the second study, type A responders were randomized to receive placebo, 5000 U or 10, 000 U of toxin.
Most patients treated with Zevalin experience a severely reduced number of blood cells, which needs some time to normalise again after treatment. Most patients treated with Zevalin experience a severe reduction in the number of blood cells. It takes some time to return to normal after treatment.
Lisonorm tablet contains 5mg amlodipine/ 10mg lisinopril and is an anti-hypertensive indicated in patients with blood pressure adequately controlled with lisinopril and amlodipine given concurrently at the same dose level. Lisonorm tablet contains 5 mg of amlodipine and 10 mg of lisinopril. It is indicated for the treatment of hypertension in patients whose blood pressure is adequately controlled with lisinopril and amlodipine.
@@
She walked out on their wedding , and she currently is single and holds intense anger towards Richard for ruining their wedding, and also became furious to learn she is pregnant, even though Richard told her he got a vasectomy. She walked out on their wedding. She currently is single. She holds intense anger towards Richard for ruining their wedding. She became furious to learn she is pregnant. Richard told her he got a vasectomy.
Southfort is an unincorporated community in DeSoto County, Florida, United States, located approximately south of Fort Ogden on U.S. Route 17. Southfort is an unincorporated community in DeSoto County, Florida. It is located approximately south of Fort Ogden on U.S. Route 17.
Parke played for powerhouse youth club FC Delco, and played four years of college soccer at Drexel University, before being drafted by the MetroStars with 60th and last overall pick of the 2004 MLS SuperDraft. Parke played for powerhouse youth club FC Delco. He played four years of college soccer at Drexel University. He was drafted by the MetroStars with 60th and last overall pick of the 2004 MLS SuperDraft.
Auguste Le Duc started on a bridge between the two islands, which was quickly swept away by the forces of nature. Auguste Le Duc started on a bridge between the two islands. The bridge was quickly swept away by the forces of nature.
@@
Appendix C Extended Prompting Variants
Few-shot prompts.
As noted in Section 4.2, we explored the use of larger few-shot sampling on the instruction-tuned models, using 10 in-context learning examples instead of the 5 selected for our core experiments. Results of these additional experiments are presented in Table 10 for the best-performing instruction-tuned Alpaca 7B model.
Prompting with 10 examples led to relatively minor gains in SARI in most cases, with a marked tendency to delete more content in the rephrased version. BertScore and BLEU were virtually similar. Extending the in-context samples also resulted in CPL losses with domain shift, up to 10 points with 10-shot–BI on the DeSSE test set, notably, although the reverse was observed with 10-shot-DE on the BiSECT test, with a 2.59 points gain. A minor CPL loss was also observed with 10-shot-BI over BiSECT.
Domain shift losses were not totally surprising considering the in-context conditioning on samples from the very different DeSSE and BiSECT domains. Considering these results, in particular CPL losses, we selected 5-shot as our primary method for few-shot in-context learning.
Zero-shot prompts.
The zero-shot prompt selected for our experiments was chosen after preliminary testing of different manually engineered variants. No clear better prompting candidate emerged from this initial phase, although it is always possible that an alternative prompt could lead to better result with a given model. To provide further information on zero-shot prompting variants, in Table 11 we report additional results with the following prompts, against GPT-4 Turbo (gpt-4-1106-preview, accessed February 2024):
  • Prompt O: "Split the sentence below into separate sentences."
  • Prompt A:"Split the following sentence into multiple shorter sentences, while preserving the original meaning as closely as possible."
  • Prompt B:"Rephrase the given sentence to make it more concise and direct, without changing the core meaning."
  • Prompt C:"Split the following sentence into multiple concise sentences, separating each distinct idea."
Prompt O is our original prompt, used throughout our experiments, and prompts A-C are alternative manually engineered prompts for the task. [^13] We thank the anonymous ARR reviewer who suggested these specific prompt variants as additional test cases. Leaving aside prompt B, which resulted in almost no compliance and severe degradation overall, the other two prompt variants achieved relatively similar results compared to Prompt O. Overall, the latter still achieved better results on all metrics on DeSSE, and all but SARI on BiSECT. All prompt variants were outperformed by the best fine-tuned 6.9B variants, except on compliance.
Test Method BESC BLEU SARI add keep del CPL
DeSSE 5-shot-DE 92.51 74.90 56.08 27.76 88.92 51.56 68.65
10-shot-DE 92.75 75.10 58.07 31.49 88.46 54.26 71.83
5-shot-BI 88.04 51.00 42.53 11.51 73.53 42.55 31.94
10-shot-BI 87.40 47.90 40.09 8.55 71.44 40.28 21.63
BiSECT 5-shot-BI 86.99 36.80 35.82 8.57 60.46 38.43 65.86
10-shot-BI 86.60 35.70 38.06 10.08 60.03 44.09 64.31
5-shot-DE 87.63 38.50 28.76 5.46 59.92 20.89 66.38
10-shot-DE 87.73 38.00 29.32 5.74 59.62 22.60 68.97
Table 10: Extended few-shot results with the Alpaca-7B model using examples from DeSSE (DE) and BiSECT (BI).
Method Test BESC BLEU SARI add keep del CPL
GPT-4 Prompt O DeSSE 84.59 70.90 63.69 35.24 89.76 66.09 99.20
GPT-4 Prompt A DeSSE 79.92 59.40 56.52 25.74 82.94 60.86 99.80
GPT-4 Prompt B DeSSE 61.87 15.40 26.27 2.58 37.18 39.05 1.38
GPT-4 Prompt C DeSSE 79.27 58.50 55.45 24.24 81.89 60.24 100.00
pythia-6.9B-DE (6K) DeSSE 88.78 81.80 69.47 45.75 92.26 70.39 94.64
GPT-4 Prompt O BiSECT 72.04 34.70 37.18 9.65 60.33 41.55 100.00
GPT-4 Prompt A BiSECT 67.78 29.30 42.54 10.36 59.92 57.33 100.00
GPT-4 Prompt B BiSECT 65.37 18.40 43.38 9.04 49.90 71.20 3.96
GPT-4 Prompt C BiSECT 68.14 30.60 42.59 10.46 60.61 56.69 100.00
pythia-6.9B-BI (30K) BiSECT 76.32 42.40 45.06 16.81 65.93 52.45 95.17
Table 11: Zero-shot prompt variant results on DeSSE and BiSECT with GPT4. Pythia-6.9B results are provided as a reference. Best results amongst GPT-4 variants are shown in bold.
Appendix D Evaluation Metrics on Wiki-BM
@@
In section 7, we used portions of the Wiki-BM corpus (Zhang et al., 2020a) for our qualitative evaluation of LLM-based SPRP output. In Table 12, we provide the results in terms of metrics for the complete Wiki-BM test set, along with results from AlAjlouni and Li (2023), as representative of an alternative method based on fine-tuning an encoder-decoder pretrained T5 model, and GPT4 Turbo, as one of the largest LLMs to date. The caveats mentioned in Section 6 regarding the latter apply to these results as well.
@@
The results for our LLM-based approach are in line with those obtained on the DeSSE test set across metrics: slightly lower on BESC, BLEU, and SARI, but higher in terms of compliance (CPL). The Pythia-6.9B model fine-tuned with LoRA notably outperformed the reported scores in AlAjlouni and Li (2023) by 13.4 and 5.1 points on the BLEU and SARI metrics, respectively. GPT-4 Turbo achieved the best results on all metrics, but by relatively small margins overall. These results confirm the viability of relatively smaller LLMs, tuned with LoRA on small amounts of data, for the SPRP task.
@@
pythia-6.9B-DE 72.90 65.10 36.14 91.15 68.10 85.50 98.19
TMWS AlAjlouni and Li (2023) 59.50 60.00 - - - - -
GPT-4 Turbo 74.40 68.63 39.96 91.90 74.03 86.23 99.70
Table 12: Metrics results on the Wiki-BM benchmark test set. Comparison between Pythia-6.9B fine-tuned on DeSSE data, results reported in (AlAjlouni and Li, 2023), and GPT4 Turbo (accessed on February 2024).
Appendix E 3-Way Ranking Evaluation Protocol
@@
The 3-way ranking manual evaluation involved 11 human evaluators, all researchers with at least upper-intermediate knowledge of English, who were not paid for the task. The evaluation was split into two tasks: for one task, the evaluators were asked to rank the output of the fine-tuned Pythia-6.9B model and the output of the ABCD model; for the other task, they had to rank the output of the fine-tuned Pythia-6.9B model and the output of the BiSECT model. Each of the two tasks involved 100 samples. Evaluators were also asked to ignore minor capitalisation and tokenisation issues, which were typically present in the output of the ABCD and BiSECT models, as detected in preliminary data inspection.
@@
Appendix F Qualitative Evaluation Protocol
@@
Appendix G Initial Multilingual Results